Cartwright Brothers’ stock is currently selling for $40 a share. The stock is expected to pay a $4 dividend at the end of the year. The stock’s dividend is expected to grow at a constant rate of 7 percent a year forever. The risk-free rate (kRF) is 7 percent and the market risk premium (kM – kRF) is 5 percent. What is the stock’s beta?

a.

5.00

b.

1.00

c.

2.00

d.

3.00

e.

4.00
Answer #1

Using Gordon Growth model,

Stock Price = P0 = D1/(r - g)

where, P0 = current stock price = 40

g = growth rate = 0.07

D1 = dividend = 4

r = rate of return

=> 40 = 4/(r - 0.07) => r = 0.17 = 17%

Using the CAPM model,

r = kRF + β(kM - kRF)

given kRF = 0.07

(kM - kRF) = 0.05

=> 0.17 = 0.07 + β*0.05

=> β = 2

Hence, (c) is the correct ooption